### v1

@@ Abstract
The evolution of diffusion models has greatly impacted video generation and understanding. Particularly, text-to-video diffusion models (VDMs) have significantly facilitated the customization of input video with target appearance, motion, etc. Despite these advances, challenges persist in accurately distilling motion information from video frames. While existing works leverage the consecutive frame residual as the target motion vector [16], they inherently lack global motion context and are vulnerable to frame-wise distortions. To address this, here we present Spectral Motion Alignment (SMA), a novel framework that refines and aligns motion vectors using Fourier and wavelet transforms. SMA learns motion patterns by incorporating frequency-domain regularization, facilitating the learning of whole-frame global motion dynamics, and mitigating spatial artifacts. Extensive experiments demonstrate SMA’s efficacy in improving motion transfer while maintaining computational efficiency and compatibility across various video customization frameworks.
Keywords:
Diffusion models Video Motion transfer Wavelet transform
Figure 1: One-shot Video Motion Transfer via Spectral Motion Alignment using Cascaded Video Diffusion. SMA facilitates the capture of long-range (top) and complex (bottom) motion patterns within videos. Visit here for comprehensive view of the videos.
@@ 1 Introduction
The progression of diffusion models [31, 12] has played a pivotal role in advancing video generation and comprehension. Notably, recent advancements in text-conditioned Video Diffusion Models (VDMs, [13, 11, 30, 4, 9, 45, 3, 22]) have significantly enhanced the generation of highly realistic and superior-quality videos. Given the multifaceted nature of the video, encompassing motion dynamics, foreground object appearance, etc., several studies [49, 39, 16, 42, 41, 27, 36] aim to disentangle, modify, and compose these signals according to user intent. Additionally, various approaches specifically target to customize the motion, pose, shape, or appearance of real-world videos.
In the context of motion customization, multiple studies explore how to distill motion information from pre-trained video diffusion models. Recent research suggests that motion patterns are inherently encoded in the underlying dependencies between latent frames or epsilon noises. As noted by [49], videos with similar motion tend to exhibit similar connectivity between latent frames. Additionally, [16] utilizes residual vectors between consecutive latent frames as "motion vectors," in line with optical flow principles, where residuals in consecutive frames represent motion dynamics. Specifically, they finetune temporal attention layers of pretrained VDM to align the ground-truth image-space residuals with their predicted denoised estimates. Their work reveals that aligning these predicted and ground-truth motion vectors corresponds to aligning epsilon (noise) residuals, readily obtainable from any off-the-shelf video diffusion models.
@@
While these efforts have made technical advances in motion distillation, the pitfalls of using latent frame residuals as motion vectors have not been fully studied. This paper focuses on the potential limitations of current motion estimation in diffusion models: (a) lack of global context information, and (b) vulnerability to frame-wise distortions. Since frame residuals may capture local motion patterns but are blind to whole-frame motion dynamics, for better motion dynamics modeling, we have to understand the whole-frame global context information during motion distillation. Furthermore, while the frame residuals contain motion information, they may also contain inevitable disruptive variations that are unrelated to motion. These variations may include abrupt changes in the background, distortions, consecutive frame inconsistencies, or lighting condition changes.
To address these challenges, here we present Spectral Motion Alignment (SMA), a novel Fourier and wavelet-transform-based motion vector refinement and alignment framework. This includes two primary components: First, to learn the global motion context, we propose the spectral alignment loss between predicted and ground-truth motion vectors within the wavelet domain. This facilitates the learning of multi-scale motion dynamics by leveraging rich wavelet-domain representations of video considering the global frame transitions. Moreover, to mitigate the spatial artifacts and inconsistency in motion vectors, we propose 2D FFT-based motion vector refinement that aligns the amplitude and phase spectrum of ground truth and predicted motion vectors with prioritizing low-frequency components. This is because the high-frequency components in motion vectors may be associated with frame-wise non-motion-related artifacts (Figure 6). To sum up, we encourage accurate motion transfer via harmonized global and local levels of frequency-domain regularization.
Our contributions are summarized as follows:
  • We introduce the Spectral Motion Alignment (SMA), a frequency-domain motion alignment framework that learns the underlying motion dynamics of input video via frequency-based regularization. Since conventional methods mostly rely on pixel-domain motion distillation, the proposed framework is orthogonal to existing motion customization models.
  • SMA imposes negligible memory and computational burdens, as most off-the-shelf VDMs can readily compute estimates of motion vectors. For instance, VMC [16] with SMA demonstrates lightweight (15GB vRAM) and rapid (< 5 minutes) training.
  • We validate the advantage of SMA across diverse motion patterns and subjects, and across various video motion transfer frameworks such as Video Diffusion-based [49], Cascaded Video Diffusion-based [16], Text-to-Image Diffusion-based [38], and ControlNet-based models [7].
@@ 2 Preliminaries
Diffusion models [31, 12] generate samples from the Gaussian noise through reverse denoising processes. Given a clean sample ${\boldsymbol{x}}_{0}\sim p_{\operatorname{data}}({\boldsymbol{x}})$, the forward process is defined as a Markov chain with forward conditional densities
$\begin{split}p({\boldsymbol{x}}_{t}\>|\>{\boldsymbol{x}}_{t-1})={\mathcal{N}}({\boldsymbol{x}}_{t}\>|\>\beta_{t}{\boldsymbol{x}}_{t-1},(1-\beta_{t})I)\\ p_{t}({\boldsymbol{x}}_{t}\>|\>{\boldsymbol{x}}_{0})={\mathcal{N}}({\boldsymbol{x}}_{t}\>|\>\sqrt{{\bar{\alpha}}}{\boldsymbol{x}}_{0},(1-{\bar{\alpha}})I),\end{split}$ (1)
where ${\boldsymbol{x}}_{t}\in\mathbb{R}^{d}$ is a noisy latent variable at time $t$ that has the same dimension as ${\boldsymbol{x}}_{0}$, and $\beta_{t}$ denotes an increasing sequence of noise schedule where $\alpha_{t}\coloneqq 1-\beta_{t}$ and ${\bar{\alpha}}_{t}\coloneqq\Pi_{i=1}^{t}\alpha_{i}$. Then, the goal of diffusion model training is to obtain a residual denoiser ${\boldsymbol{\epsilon}}_{\theta^{*}}$:
$\theta^{*}:=\mathop{\mathrm{argmin}}_{\theta}{\mathbb{E}}_{{\boldsymbol{x}}_{t}\sim p_{t}({\boldsymbol{x}}_{t}\>|\>{\boldsymbol{x}}_{0}),{\boldsymbol{x}}_{0}\sim p_{\operatorname{data}}({\boldsymbol{x}}_{0}),{\boldsymbol{\epsilon}}\sim{\mathcal{N}}(0,I)}\big{[}\left\lVert{\boldsymbol{\epsilon}}_{\theta}({\boldsymbol{x}}_{t},t)-{\boldsymbol{\epsilon}}\right\rVert\big{]}.$ (2)
@@
where ${\boldsymbol{\epsilon}}\sim{\mathcal{N}}(0,{\boldsymbol{I}})$ and $\tilde{\beta}_{t}\coloneqq\frac{1-{\bar{\alpha}}_{t-1}}{1-{\bar{\alpha}}_{t}}\beta_{t}$. To accelerate sampling, DDIM [32] further proposes another sampling method as follows:
${\boldsymbol{x}}_{t-1}=\sqrt{{\bar{\alpha}}_{t-1}}\hat{{\boldsymbol{x}}}_{0}(t)+\sqrt{1-{\bar{\alpha}}_{t-1}-\eta^{2}\tilde{\beta_{t}}^{2}}{\boldsymbol{\epsilon}}_{\theta^{*}}({\boldsymbol{x}}_{t},t)+\eta\tilde{\beta}_{t}{\boldsymbol{\epsilon}},$ (4)
where $\eta\in[0,1]$ is a stochasticity parameter, and $\hat{{\boldsymbol{x}}}_{0}(t)$ is the denoised estimate which can be equivalently derived using Tweedie’s formula [8]:
@@
2.1.1 Video Diffusion Models.
Video diffusion models [13, 11, 45] further attempt to model the video data distribution. Specifically, Let $({\boldsymbol{v}}^{n})_{n\in\{1,\dots,N\}}$ represents the $N$-frame input video sequence. Then, for a given $n$-th frame ${\boldsymbol{v}}^{n}\in\mathbb{R}^{d}$, let ${\boldsymbol{v}}^{1:N}\in\mathbb{R}^{N\times d}$ represents a whole video vector. Let $\boldsymbol{v}_{t}^{n}=\sqrt{\bar{\alpha}_{t}}\boldsymbol{v}^{n}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon}_{t}^{n}$ represents the $n$-th noisy frame latent sampled from $p_{t}(\boldsymbol{v}_{t}^{n}|\boldsymbol{v}^{n})$, where $\boldsymbol{\epsilon}_{t}^{n}\sim\mathcal{N}(0,I)$. We similarly define $({\boldsymbol{v}}_{t}^{n})_{n\in 1,\dots,N}$, ${\boldsymbol{v}}_{t}^{1:N}$, and ${\boldsymbol{\epsilon}}^{1:N}$. The goal of video diffusion model training is then to obtain a residual denoiser ${\boldsymbol{\epsilon}}_{\theta}$ with textual condition $c$ and video input that satisfies:
@@ 2.2 Fourier and Wavelet Analysis
Spectral analysis techniques transform time-domain or pixel-domain signals (such as video frames) into the frequency domain, revealing the frequency components and their intensities. This transformation is valuable in video understanding as it helps in identifying repetitive motion patterns and underlying structures that may not be visible in the time or pixel domain.
@@
$|\mathcal{F}_{{\boldsymbol{v}}^{n}}(a,b)|=\sqrt{R(a,b)^{2}+I(a,b)^{2}},\angle\mathcal{F}_{{\boldsymbol{v}}^{n}}(a,b)=\arctan\Big{(}\frac{I(a,b)}{R(a,b)}\Big{)}.$ (8)
Here, amplitude often indicates the existence of corresponding spatial frequencies within the image, often characterized by edges, textures, patterns, etc. Phase often encodes the structural spatial relationship.
Wavelet Transform. We first provide a brief introduction to the frame theory. Frames in signal processing are an extension of basis sets but with a relaxation that allows for redundancy. Let $\Psi=[\psi_{1}\dots\psi_{m}]\in\mathbb{R}^{n\times m}$, where $\{\psi_{k}\}_{k=1}^{m}$ is a set of functions in a Hilbert space $H$. Then, $\{\psi_{k}\}_{k=1}^{m}$ is called as a frame if it satisfies the following inequality:
$\alpha\|f\|^{2}\leq\|\langle f,\Psi\rangle\|^{2}\leq\beta\|f\|^{2},\forall f\in H$ (9)
where $c_{k}\coloneqq\langle f,\psi_{k}\rangle$ is a $k$-th frame expansion coefficient, and $\alpha,\beta$ are frame bounds. The original signal $f$ can be exactly reconstructed from expansion coefficients using the dual frame operator $\tilde{\Psi}$ which satisfies the so-called frame condition: $\tilde{\Psi}\Psi^{T}=I$.
Wavelet frames, renowned for capturing multi-resolution scale features, are among the most prevalently utilized frame representations in signal processing. Let $\psi(t)$ represent a mother wavelet that can be shifted and scaled. For a function ${\boldsymbol{v}}(t)\in L^{2}(\mathbb{R})$, the wavelet transform can be expressed as:
@@
Then, the original signal can be recovered from inverse DWT. In practice, this discrete wavelet transform can be implemented by convolution using an appropriate choice of filter bank.
@@ 3 Spectral Motion Alignment
Given an input video, our main goal is to develop a novel frequency-domain motion alignment framework that (a) distills the motion patterns $M^{*}$ of input video, and (b) transfers the motion patterns $M^{*}$ to output video within varied contexts, e.g. Cars with motion $M^{*}$rightarrow Tanks with motion $M^{*}$.
Conventional motion distillation methods mostly rely on pixel-domain motion distillation objectives. However, frequency-domain analysis can further capture underlying motion patterns across a spectrum of frequency levels that mainly constitute motion. Accordingly, we propose Spectral Motion Alignment (SMA), a novel Fourier and wavelet-transform-based motion vector refinement and alignment framework. More details follow.
@@ 3.1 Denoised Motion Vector Estimation
To distill the motion information, we first estimate the motion vector of the input video following [16]. The intuition is that residual vectors between consecutive frames may include information about the motion trajectories. Define the $n$-th frame residual vector, namely motion vector at time $t\geq 0$ as
@@
where the epsilon residual vector $\delta{\boldsymbol{\epsilon}}_{t}^{n}$ is similarly defined. This $\delta{\boldsymbol{v}}_{t}^{n}$ can be acquired through the following diffusion kernel [16]:
@@
Moreover, [16] shows that the ground-truth motion vector in pixel space $\delta{\boldsymbol{v}}_{0}^{n}$ can be derived as follows:
@@
Similarly, one can obtain the denoised motion vector estimate $\delta\hat{{\boldsymbol{v}}}_{0}^{n}$ by using Tweedie’s formula as follows:
@@
In the context of motion transfer, the authors in [16] finetune the temporal attention layers of VDM by aligning each motion vector $\delta{\boldsymbol{v}}_{0}^{n}$ and its denoised estimate $\delta\hat{{\boldsymbol{v}}}_{0}^{n}(t)$:
@@
While these advancements in motion distillation mark significant progress, Figure 1, 4 and 6 indicate that [16] still has potential for further refinement.
@@ 3.2 Spectral Global Motion Alignment
One of the primary limitations in (17) is that it may not fully encapsulate the global motion dynamics. This is partly because [16] employs ‘frame-wise’ matching losses, such as $\ell_{2}$-distance. While this is effective, it may overlook the comprehensive motion dynamics by focusing solely on pairwise frame comparisons.
@@
$\ell_{\text{global}}(\delta{\boldsymbol{v}}_{0},\delta\hat{{\boldsymbol{v}}}_{0}(t))=\mathbb{E}_{t,s,j,k}\Big{[}\|{\mathcal{W}}_{\delta{\boldsymbol{v}}_{0,s}}(j,k)-{\mathcal{W}}_{\delta\hat{{\boldsymbol{v}}}_{0,s}(t)}(j,k)\|_{1}\Big{]}.$ (19)

### v2

@@ Abstract
Diffusion models have significantly facilitated the customization of input video with target appearance while maintaining its motion patterns. To distill the motion information from video frames, existing works often estimate motion representations as frame difference or correlation in pixel-/feature-space. Despite its simplicity, these methods have unexplored limitations, including lack of understanding of global motion context, and the introduction of motion-independent spatial distortions. To address this, we present Spectral Motion Alignment (SMA), a novel framework that refines and aligns motion representations in the spectral domain. Specifically, SMA learns spectral motion representations, facilitating the learning of whole-frame global motion dynamics, and effectively mitigating motion-independent artifacts. Extensive experiments demonstrate SMA’s efficacy in improving motion transfer while maintaining computational efficiency and compatibility across various video customization frameworks.
Figure 1: One-shot Video Motion Transfer via Spectral Motion Alignment using Cascaded Video Diffusion Models. SMA facilitates the capture of long-range (left) and complex (right) motion patterns within videos. Visit https://geonyeong-park.github.io/spectral-motion-alignment/ for a comprehensive view of the videos.
@@ 1 Introduction
Given the multifaceted nature of the video, encompassing motion dynamics, appearance, etc., several studies aim to disentangle and control these signals according to user intent. Recently, diffusion models (Sohl-Dickstein et al. 2015; Ho, Jain, and Abbeel 2020) has played a pivotal role in video customization, owing to their superior sampling ability.
In the context of motion customization using diffusion models, our goal is to transfer the motion patterns from an input video to the customized output video. This necessitates the accurate estimation and extraction of motion information from the input video. While fundamental techniques such as optical flow are effective for motion estimation, integrating these into diffusion models for customization is nontrivial.
To address these challenges, recent researches suggest that motion patterns are inherently encoded in the underlying dependencies between frames or epsilon noises. For example, (Zhao et al. 2023b) have observed that videos with similar motion tend to exhibit similar connectivity between latent frames. Additionally, (Jeong, Park, and Ye 2023) utilizes residual vectors between consecutive frames as "motion vectors," in line with optical flow principles, assuming frame residuals represent motion dynamics. Specifically, they finetune pretrained VDM to align the ground-truth pixel-space residuals with their predicted denoised estimates. Thus, these works leverage the pixel-space differences between input frames as a proxy of motion reference.
While these motion representations can be obtained from off-the-shelf video diffusion models efficiently, current simple approximations have several adverse impacts. First, they may fail to capture the global context of motion. Since frame residuals may capture local motion patterns but are blind to whole-frame motion dynamics, for better motion dynamics modeling, we have to understand the whole-frame global context information during motion distillation. Furthermore, while the pixel- or feature-space residuals contain motion information, they may also contain inevitable disruptive variations that are unrelated to motion. These variations may include abrupt changes in the background, lightning, or other frame inconsistencies, leading to less reliable representation.
To address these challenges, we introduce Spectral Motion Alignment (SMA), a novel framework for refining and aligning motion representations in the spectral domain, based on intuition that the motion may be well represented by its inherent frequency components. This framework includes two primary components: First, to capture the global motion context, we propose a spectral alignment loss between predicted and ground-truth motion vectors within the wavelet domain. This facilitates the learning of multi-scale motion dynamics by leveraging rich wavelet-domain representations of video considering the global frame transitions. Moreover, to mitigate the spatial artifacts and inconsistency in motion vectors, we propose 2D FFT-based motion vector refinement that aligns the amplitude and phase spectrum of ground truth and predicted motion vectors with prioritizing low-frequency components. This is because the high-frequency components in motion representations may be associated with frame-wise motion-independent artifacts (Figure 6). In summary, we encourage accurate motion transfer via harmonized global and local levels of spectral domain alignment. Our contributions are summarized as follows:
@@
  • We introduce the Spectral Motion Alignment (SMA), a frequency-domain motion alignment framework that learns the underlying motion dynamics of input video via frequency-based regularization. SMA is orthogonal and compatible to most of existing motion customization models as they often only rely on either pixel or feature space representations.
  • SMA imposes negligible memory and computational burdens, as most off-the-shelf VDMs can readily compute motion vectors estimates. For instance, VMC (Jeong, Park, and Ye 2023) with SMA demonstrates lightweight (15GB vRAM) and rapid ($<$ 5 min) training.
  • We validate the efficacy of SMA across diverse motion patterns, subjects, and various video motion transfer frameworks including Video Diffusion-based (Zhao et al. 2023b), Cascaded Video Diffusion-based (Jeong, Park, and Ye 2023), T2I Diffusion-based (Wu et al. 2023), and ControlNet-based models (Chen et al. 2023b).
@@ 2 Preliminaries
Diffusion models (Sohl-Dickstein et al. 2015; Ho, Jain, and Abbeel 2020) generate samples from the Gaussian noise through reverse denoising processes. We denote a clean sample ${\boldsymbol{x}}_{0}\sim p_{\operatorname{data}}({\boldsymbol{x}})$, a noisy latent ${\boldsymbol{x}}_{t}\in\mathbb{R}^{d}$ at time $t$, $\beta_{t}$ as an increasing sequence of noise schedule, $\alpha_{t}\coloneqq 1-\beta_{t}$, and ${\bar{\alpha}}_{t}\coloneqq\Pi_{i=1}^{t}\alpha_{i}$. Then the goal of diffusion model training is to optimize a denoiser ${\boldsymbol{\epsilon}}_{\theta^{*}}$:
$\theta^{*}:=\mathop{\mathrm{argmin}}_{\theta}{\mathbb{E}}_{{\boldsymbol{x}}_{t},{\boldsymbol{x}}_{0},{\boldsymbol{\epsilon}}}\big{[}\left\lVert{\boldsymbol{\epsilon}}_{\theta}({\boldsymbol{x}}_{t},t)-{\boldsymbol{\epsilon}}\right\rVert\big{]}.$ (1)
@@
where ${\boldsymbol{\epsilon}}\sim{\mathcal{N}}(0,{\boldsymbol{I}})$ and $\tilde{\beta}_{t}\coloneqq\frac{1-{\bar{\alpha}}_{t-1}}{1-{\bar{\alpha}}_{t}}\beta_{t}$. To accelerate sampling, DDIM (Song, Meng, and Ermon 2020) further proposes another sampling method as follows:
${\boldsymbol{x}}_{t-1}=\sqrt{{\bar{\alpha}}_{t-1}}\hat{{\boldsymbol{x}}}_{0}(t)+\sqrt{1-{\bar{\alpha}}_{t-1}-\eta^{2}\tilde{\beta_{t}}^{2}}{\boldsymbol{\epsilon}}_{\theta^{*}}({\boldsymbol{x}}_{t},t)+\eta\tilde{\beta}_{t}{\boldsymbol{\epsilon}},$
where $\eta\in[0,1]$ controls stochasticity, and $\hat{{\boldsymbol{x}}}_{0}(t)$ is the denoised estimate which can be equivalently derived using Tweedie’s formula (Efron 2011):
@@
Video Diffusion Models.
Video diffusion models (Ho et al. 2022b, a; Zhang et al. 2023a) further attempt to model the video data distribution. Specifically, Let $({\boldsymbol{v}}^{n})_{n\in\{1,\dots,N\}}$ represents the $N$-frame input video sequence. Then, for a given $n$-th frame ${\boldsymbol{v}}^{n}\in\mathbb{R}^{d}$, let ${\boldsymbol{v}}^{1:N}\in\mathbb{R}^{N\times d}$ represents a whole video vector. Let $\boldsymbol{v}_{t}^{n}=\sqrt{\bar{\alpha}_{t}}\boldsymbol{v}^{n}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon}_{t}^{n}$ represents the $n$-th noisy frame latent sampled from $p_{t}(\boldsymbol{v}_{t}^{n}|\boldsymbol{v}^{n})$, where $\boldsymbol{\epsilon}_{t}^{n}\sim\mathcal{N}(0,I)$. We similarly define $({\boldsymbol{v}}_{t}^{n})_{n\in 1,\dots,N}$, ${\boldsymbol{v}}_{t}^{1:N}$, and ${\boldsymbol{\epsilon}}^{1:N}$. The goal of video diffusion model training is then to obtain a residual denoiser ${\boldsymbol{\epsilon}}_{\theta}$ with textual condition $c$ and video input that satisfies:
@@ 2.2 Fourier and Wavelet Analysis
Spectral analysis techniques transform time-domain or pixel-domain signals (such as video frames) into the frequency domain, revealing the frequency components and their intensities.
@@
$\begin{split}|\mathcal{F}_{{\boldsymbol{v}}^{n}}(a,b)|=\sqrt{R(a,b)^{2}+I(a,b)^{2}},\\ \angle\mathcal{F}_{{\boldsymbol{v}}^{n}}(a,b)=\arctan\Big{(}\frac{I(a,b)}{R(a,b)}\Big{)}.\end{split}$ (6)
Wavelet Transform. Wavelet frames, renowned for capturing multi-resolution scale features, are among the most prevalently utilized frame representations in signal processing. Let $\psi(t)$ represent a mother wavelet that can be shifted and scaled. For a function ${\boldsymbol{v}}(t)\in L^{2}(\mathbb{R})$, the wavelet transform can be expressed as:
@@
The original signal can be recovered from inverse DWT. In practice, this discrete wavelet transform can be implemented by convolution using an appropriate choice of filter bank.
@@ 3 Spectral Motion Alignment
Our main goal is to develop a novel spectral domain motion alignment framework that capture underlying complex motion patterns across a spectrum of frequency levels that mainly constitute motion. This is valuable in video understanding and customization as it helps in identifying repetitive motion patterns and underlying structures that may not be visible in the time or pixel domain. It is in orthogonal (and compatible) with conventional methods based on pixel- or feature-domain motion representations.
@@ 3.1 Denoised Motion Vector Estimation
To distill the motion information, we first estimate the initial motion representations (Jeong, Park, and Ye 2023; Zhao et al. 2023b) in pixel space. For this, we follow VMC (Jeong, Park, and Ye 2023) as an representative example. The intuition is that residual vectors between consecutive frames may include information about the motion trajectories. Define the $n$-th frame residual vector, namely motion vector at time $t\geq 0$ as
@@
where the epsilon residual vector $\delta{\boldsymbol{\epsilon}}_{t}^{n}$ is similarly defined. This $\delta{\boldsymbol{v}}_{t}^{n}$ can be acquired through the following diffusion kernel:
@@
Given that, the ground-truth motion vector in pixel space $\delta{\boldsymbol{v}}_{0}^{n}$ can be derived as follows:
@@
Similarly, one can obtain the denoised estimate version of these motion representations $\delta\hat{{\boldsymbol{v}}}_{0}^{n}$ by using Tweedie’s formula as follows:
@@
In the context of motion transfer, we aim to align the ground-truth and estimated motion vectors by fine-tuning the pre-trained VDM (Jeong, Park, and Ye 2023; Zhao et al. 2023b):
@@
While these advancements in motion distillation mark significant progress, Figure 1, 4 and 6 indicate that existing methods still has potential for further refinement.
@@ 3.2 Spectral Global Motion Alignment
One of the primary limitations in (14) is that it may not fully encapsulate the global motion dynamics. Specifically, it locally focuses on pairwise frame comparisons which may lead to overlooking the comprehensive motion dynamics of given object overall frames.
@@
$\begin{split}\ell_{\text{global}}&(\delta{\boldsymbol{v}}_{0},\delta\hat{{\boldsymbol{v}}}_{0}(t))=\\ &\mathbb{E}_{t,s,j,k}\Big{[}\|{\mathcal{W}}_{\delta{\boldsymbol{v}}_{0,s}}(j,k)-{\mathcal{W}}_{\delta\hat{{\boldsymbol{v}}}_{0,s}(t)}(j,k)\|_{1}\Big{]}.\end{split}$ (16)